ROTATIONAL AND HYPERFINE STRUCTURE IN THE B̃ - X̃ TRANSITION OF NIOBIUM METHYLI-DYNE, NbCH

D. J. CLOUTHIER, Department of Chemistry, University of Kentucky, Lexington, KY 40506-0055; \underline{C. T. KINGSTON} and A. J. MERER, Department of Chemistry, University of British Columbia, 2036 Main Mall, Vancouver, B.C. V6T 1Z1, Canada.

High resolution spectra have been taken for some of the components of the perturbed (0,0) band of the B̃^{3}Δ_{1} – X̃^{3}Δ_{1} system of NbCH near 600 nm. Extensive Nb hyperfine structure (I = 9/2) is found in all the rotational lines, consistent with the ground state electron configuration (5sσ)^{1} (4dδ)^{1}. Experiments with NbCD are planned, and it is hoped to report the hyperfine constants and the molecular structure at the meeting.